### v1

# OraPO: Oracle-educated Reinforcement Learning for Data-efficient and Factual Radiology Report Generation

Abstract
Radiology report generation (RRG) aims to automatically produce clinically faithful reports from chest X-ray images. Prevailing work typically follows a scale-driven paradigm, by multi-stage training over large paired corpora and oversized backbones, making pipelines highly data- and compute-intensive. In this paper, we propose Oracle-educated GRPO (OraPO) with a FactScore-based reward (FactS) to tackle the RRG task under constrained budgets. OraPO enables single-stage, RL-only training by converting failed GRPO explorations on rare or difficult studies into direct preference supervision via a lightweight oracle step. FactS grounds learning in diagnostic evidence by extracting atomic clinical facts and checking entailment against ground-truth labels, yielding dense, interpretable sentence-level rewards. Together, OraPO and FactS create a compact and powerful framework that significantly improves learning efficiency on clinically challenging cases, setting the new SOTA performance on the CheXpert Plus dataset (0.341 in F1) with 2–3 orders of magnitude less training data using a small base VLM on modest hardware.
Figure 1: Illustration of mainstream data/compute-intensive pipelines (upper-left) versus our data-efficient pipeline (upper-right). Bottom: our method achieves the SOTA performance for RRG with less than 0.1% of the training samples (vs. 1.27 M) used by best-performing baselines and a much smaller model, demonstrating strong performance under tight data and compute budgets.
footnotetext: Work done during an internship at Oracle Health & AI. Joint senior authors.
1 Introduction
Radiology report generation (RRG) from chest X-rays aims to turn imaging evidence into a clinically usable narrative. This is a complex multimodal task that couples disease prediction with free-text report generation, linking detected visual findings (e.g. edema, pleural effusion, and pneumothorax) to clear sentences that describe their presence, details, location, and severity. Automating this task matters in practice: the large imaging volumes and backlogs place serious strains on healthcare services and delay patient care, while radiologist staffing remains insufficient. For example, England reports a 29% radiology consultant shortfall alongside sustained vacancies [48], while in the United States there are about 14,000 openings but only 1,150 resident graduates annually [53]. In this context, AI-assisted drafting has begun to show measurable efficiency gains. A recent study [1] reported that using draft reports generated by large language models (LLMs) reduced reporting time by 24%. Given its practical importance, automated radiology report generation has rapidly become an active research area aimed at easing radiologists’ workload and streamlining workflows in hospitals and clinics.
Mainstream RRG methods typically rely on supervised fine-tuning (SFT) via multi-stage training on large corpora and/or scaling up backbone size (Fig. OraPO: Oracle-educated Reinforcement Learning for Data-efficient and Factual Radiology Report Generation, upper-left). The multi-stage training paradigm usually follows three phases: domain-specific pre-training, image–text alignment, and task-specific fine-tuning, where the first two are usually trained on combinations of large-scale, multi-source datasets such as medical textbooks, real radiology reports including but not limited to MIMIC-CXR [29], IU X-ray [17] and CheXpert Plus [8], often for dozens of epochs [62, 64]. Notably, both CheXpert Plus and MIMIC-CXR include training splits exceeding 200K image-report pairs. In addition, there is a recent trend of adopting large vision-language models (VLMs) with >>13B parameters for SFT, demanding meticulous data curation and substantial GPU budgets [32, 30].
Recently, Group Relative Policy Optimisation (GRPO) has emerged as a strong and compute-efficient RL-based learning algorithm for many hard problems with verifiable rewards. Its efficiency comes from two fronts: (i) lower memory/compute by dispensing with the trainable value critic and instead computing group-normalised advantages from the relative rewards of multiple samples [45, 31]; and (ii) a simplified training recipe that enables “RL-only” regimes (i.e., without costly SFT/alignment), avoiding massive pre-training/alignment corpora [54, 16, 10].
However, applying GRPO to the RRG task reveals two challenges. Challenge (i): we observe that vanilla GRPO often produces all-zero reward groups early in training (e.g. 30% in the first 50 steps as shown in Fig. 2 left), causing vanishing gradients and wasted rollouts. Recent variants attempt to fix this by resampling until a non-zero reward appears [75] or by enlarging the group size [55], both of which are still suboptimal due to increased cost of more rollouts. Another line interleaves SFT with RL to maintain gradient flow, yet still leaves low-quality rollouts unused, wasting costly sampling [40, 11, 76, 39, 37]. Challenge (ii): vanilla GRPO requires readily verifiable scalar rewards. Unlike multiple-choice QA, maths, or coding, where the reward is typically a binary final-answer check, radiology reports are long-form, multi-fact, and must be consistent across sentences, making reward design both critical and difficult. In practice, RLHF-style RRG work often uses reference-overlap metrics that mainly capture fluency/surface similarity (e.g. BLEU/CIDEr [33, 47]) or a report-level clinical metric [80, 44, 74]. These proxies under-penalise sentence-level factual errors and cross-sentence contradictions, yielding fluent yet clinically misleading narratives (missed positives, unsupported claims). Hence, a reward must be able to assess factual correctness of each statement of a radiology report in an efficient manner.
In this paper, we propose Oracle-educated Group Relative Policy Optimisation: OraPO, a novel RL-based learning algorithm that tackles RRG’s challenges in a data- and compute-efficient manner. OraPO addresses Challenge (i) of GRPO by dynamically injecting oracle supervision (i.e. ground-truth) for groups with all-zero rewards and teaching the policy to avoid repeating low-quality generations. Specifically, When a sampled group yields all-zero rewards (group 1 in Fig. OraPO: Oracle-educated Reinforcement Learning for Data-efficient and Factual Radiology Report Generation, upper-right), OraPO triggers a lightweight Direct Preference Optimisation (DPO) update that directly prefers the ground-truth report over the zero-reward rollouts as ready-made negatives, without extra compute or annotations. This converts failed exploration on difficult cases (hard or low-prevalence studies) into useful gradients, moving the policy away from low-quality modes and toward the ground-truth. We also introduce a simple adaptive mixing weight to increase DPO’s influence when zero-reward groups are frequent, which gradually tapers once GRPO provides informative advantages, making OraPO salvage otherwise wasted batches, stabilise training, and accelerate convergence.
Moreover, to provide OraPO with clinically faithful supervision, we introduce a FactScore-based reward [42] designed to address Challenge (ii). We treat the generated report as its own rationale, extract atomic clinical facts, and test their entailment against the ground-truth label set. Facts that support a label contribute positively; unsupported or contradictory labels incur penalties, yielding a dense, interpretable per-label reward. This stands in sharp contrast to the original GRPO reward, which assigns reward solely to the final answer while treating the reasoning trace largely as a byproduct of generation, without ensuring it is factually grounded or consistently aligned with the answer. Our reward design ensures the generated report (serving in lieu of reasoning trace) not only produces correct cognitive reasoning trace similar to that of an experienced radiologist but also yields correct diagnostic findings, which is crucial in domains such as healthcare in which safety is of paramount importance.
In summary, our contributions are threefold:
  1. 1.
    We present OraPO, an Oracle (DPO) educated GRPO that turns failed explorations into direct preference supervision and uses an adaptive mixing schedule to emphasize DPO when GRPO’s signal vanishes, significantly improving data- and compute-efficiency. To the best of our knowledge, this is the first work that integrates direct preference learning with GRPO-based RL for efficient training.
  2. 2.
    We design a FactS Reward for RL, performing fact-level entailment against ground-truth labels, yielding dense, interpretable feedback that aligns reports with diagnostic facts while avoiding long reasoning traces.
  3. 3.
    OraPO sets the new SOTA performance for RRG, surpassing the previous best model in terms of clinical effectiveness F1 score. This is achieved while being exceptionally data-efficient, with 2–3 orders of magnitude less training data. OraPO is trained on only 1K samples. In contrast, the current best model MambaXray-L is trained on 1.27M samples and all the other models are trained on at least 223K samples. At the same time, OraPO is also compute-efficient: it is based on QWen2.5-VL-3B, a moderately-sized VLM on modest hardware (4x A10 GPUs).
Figure 2: Left: Cumulative proportion of zero-reward batches (reward batch mean = 0) vs. training step on CheXpert Plus [62]. OraPO suppresses zero-reward frequency faster than naïve GRPO. Centre/Right: Class-level F1 on the CheXpert Plus validation set [62] across checkpoints for two clinically challenging and rare classes: Pneumonia (2.70%) and Fracture (4.05%). OraPO learns earlier and maintains higher F1 than naïve GRPO.
2 Related Work
Automatic Radiology Report Generation (RRG) aims to turn imaging studies (e.g., chest X-rays) into clinically usable free-text reports. Early approaches adopted seq2seq networks (e.g., CNN–RNN and Transformer) [28, 13] for training on large paired corpora such as MIMIC-CXR [29]. Recent work advances three fronts: (i) knowledge- and rule-guided generation, which injects medical priors to curb hallucination [6] and performs disease-aware image–text alignment for trustworthy text [50]; (ii) leveraging prior studies and multi-view images to improve coverage and consistency [22, 36]; and (iii) scaling/standardizing with pre-training, releasing multi-stage pre-training recipes [62, 22]. In parallel, LLM-driven RRG explores instruction-tuned or preference-aligned training [35, 32]. Despite progress, these approaches often rely on large paired corpora and/or costly computation resources, leading to deficiency.
Group Relative Policy Optimisation (GRPO) is a critic-free, reinforcement learning objective for generative LLMs, introduced in DeepSeekMath [54]. It samples multiple completions per prompt, computes group-normalised advantages from relative rewards, and updates the policy with clipped ratios to a frozen reference; removing the critic reduces memory/compute versus PPO while preserving stability [54, 21, 45, 31]. Recent variants mainly target optimisation stability: DR.GRPO [38] mitigates length bias and skewed group rewards; CISPO [9] clips importance-sampling weights; GSPO [77] adopts sequence-level clipping for MoE; DAPO [75] resamples until a non-zero reward appears. However, these methods were not designed to improve data efficiency: most target optimisation stability while leaving data volume and training epochs essentially unchanged, and some of them often increase sampling/compute [75, 59, 55].
Direct Preference Optimisation (DPO) trains a policy from preference pairs to prefer the chosen response over the rejected one [79]. Recent variants streamline or robustify this recipe: SimPO [41] applies length-normalised mean log-probabilities with a simplified reference; Cal-DPO [67] calibrates the implicit reward to human scales; ORPO [24] drops the reference via an odds-ratio objective integrated into SFT; and KTO [29] learns from unary thumbs-up/down signals instead of pairs. In our RRG framework, DPO naturally serves as an oracle step inside OraPO to convert those failed GRPO rollouts into useful preference updates against the ground-truth, leading to data- and compute-efficient learning.
3 Methodology
In this section, we first introduce the background of Group Relative Policy Optimisation (GRPO) and Direct Preference Optimisation (DPO) (§3.1). We then identify a failure mode of GRPO in radiology report generation (RRG), when a pretrained model produces many low-quality groups (e.g., zero rewards) that yield little or no gradient. This observation motivates our proposd OraPO algorithm, which efficiently converts such failed explorations into preference supervision (§3.2). Finally, we design the FactS Reward, a fact-level entailment signal that aligns reports with diagnostic labels and provides dense, interpretable feedback for reinforcement learning (§3.3).
3.1 Preliminaries
Group Relative Policy Optimisation (GRPO). GRPO [54, 21] is a computation- and memory-efficient RL objective for aligning generative models (LLMs/VLMs), introduced for reasoning-heavy settings and now used broadly because of its efficient critic-free design and proven effectiveness. It replaces the value critic with group-normalized advantages computed among multiple completions sampled for the same input. Given a study–prompt pair $(x,p)$, the policy $\pi_{\theta}$ draws a group of $K$ sequences $\{\hat{y}_{j}\}_{j=1}^{K}\sim\pi_{\theta}(\cdot\mid x,p)$ and receives scalar rewards $r_{j}=r(x,\hat{y}_{j})$. GRPO forms a group baseline and scale,
$\bar{r}=\frac{1}{K}\sum_{j=1}^{K}r_{j},\qquad\sigma=\sqrt{\frac{1}{K}\sum_{j=1}^{K}(r_{j}-\bar{r})^{2}+\varepsilon},$ (1)
and defines the group-normalised advantage:
$A_{j}=\frac{r_{j}-\bar{r}}{\sigma}.$ (2)
Let the proximal policy optimisation (PPO) ratio be:
$\rho_{j}=\exp\!\big(\log\pi_{\theta}(\hat{y}_{j}\mid x,p)-\log\pi_{\theta_{\text{old}}}(\hat{y}_{j}\mid x,p)\big),$ (3)
where $\theta_{\text{old}}$ denotes the parameters of the behaviour policy used to sample the current group (i.e., the parameters before the update). Then the GRPO loss over a batch is defined as:
$\mathcal{L}_{\mathrm{GRPO}}$ $=-\,\mathbb{E}\Big[\min\big(\rho_{j}A_{j},\ \mathrm{clip}(\rho_{j},1-\epsilon,1+\epsilon)\,A_{j}\big)\Big]$ (4)
$\quad+\lambda_{\mathrm{KL}}\ \mathbb{E}\!\left[\mathrm{KL}\!\big(\pi_{\theta}(\cdot\mid x,p)\,\|\,\pi_{\mathrm{ref}}(\cdot\mid x,p)\big)\right],$
where $\epsilon$ is the clipping threshold, $\lambda_{\mathrm{KL}}\geq 0$ controls regularization toward a frozen reference policy $\pi_{\mathrm{ref}}$. In practice, GRPO samples $K$ completions per prompt, computes $r_{j}$, normalizes within-group to obtain $A_{j}$, and updates $\pi_{\theta}$ with the clipped-ratio objective in (4), thereby achieving critic-free policy optimisation.
Direct Preference Optimisation (DPO). DPO [79] is a widely used, preference-alignment objective for generative LLMs/VLMs. It updates the policy to prefer chosen completions over rejected ones relative to a frozen reference model. Given a preference dataset $\mathcal{P}=\{(x,p,y^{+},y^{-})\}$, where $y^{+}$ is preferred to $y^{-}$ for the same input–prompt pair $(x,p)$, DPO maximizes the log-probability margin of the preferred sample after subtracting the reference log-probabilities. Let $\pi_{\theta}$ be the trainable policy and $\pi_{\mathrm{ref}}$ the frozen reference. The expectation below is taken over preference pairs drawn from $\mathcal{P}$. With temperature $\tau>0$, define
$\Delta^{+}=\log\pi_{\theta}(y^{+}\mid x,p)\;-\;\log\pi_{\mathrm{ref}}(y^{+}\mid x,p),$ (5)
$\Delta^{-}=\log\pi_{\theta}(y^{-}\mid x,p)\;-\;\log\pi_{\mathrm{ref}}(y^{-}\mid x,p),$ (6)
and the DPO loss
$\mathcal{L}_{\mathrm{DPO}}=-\,\mathbb{E}\Big[\log\sigma\!\big(\tau\,(\Delta^{+}-\Delta^{-})\big)\Big],$ (7)
where $\sigma(\cdot)$ is the logistic sigmoid.
3.2 OraPO: Oracle-educated GRPO
Exploration Failure of GRPO in RRG. On radiology report generation, applying vanilla GRPO to a base VLM with limited radiology knowledge often fails to explore productively. The root reason is that the base model carries very limited domain knowledge about radiology. Most base VLMs are pre-trained on general vision–language corpora and tasks, for example, everyday-image captioning on MS COCO [34] and open-domain visual question answering on VQA v2 [19], rather than medical imaging and reporting. Hence, it tends to miss or misstate rare and/or clinically challenging findings (e.g., small pneumothorax, subtle interstitial edema, or linear atelectasis) when asked to produce clinical narratives. For many chest X-ray inputs, the sampled group of completions is uniformly low quality with near-zero rewards. Thus, the policy receives little learning signal to drive effective exploration.
Empirically, before any domain-specific fine-tuning, the pretrained backbone attains only $0.034$ macro-F1 (Tab. 3) on clinical effectiveness. As shown in Fig. 2 (left), in the first $50$ optimisation steps approximately $30\%$ of groups have all rewards equal to zero. Under the GRPO construction, such groups give $\bar{r}\approx 0$ and $\sigma\approx\varepsilon$, so $A_{j}=\tfrac{r_{j}-\bar{r}}{\sigma}\approx 0$ for all $j$. Consequently, the clipped-ratio term in Eq. (4) contributes essentially no gradient, and the update is dominated by the KL regularizer to $\pi_{\mathrm{ref}}$ rather than by task reward. In effect, the optimizer wastes GPU resources by sampling $K$ completions per prompt without extracting useful signals, which both slows convergence and squanders valuable compute.
Zero-Reward Rate (ZRR). To resolve GRPO’s exploration failure on RRG, we first detect when a prompt yields little or no usable signal, then adapt learning accordingly. We introduce a Zero-Reward Rate (ZRR) that quantifies, per prompt, how severely GRPO has stalled.
For prompt $x_{i}$ with a GRPO group of $K$ completions $\{\hat{y}_{i,j}\}_{j=1}^{K}$ and rewards $r_{i,j}\in[0,1]$, define the raw zero-reward rate:
$z_{i}\;=\;\frac{1}{K}\sum_{j=1}^{K}\mathbf{1}\!\left[r_{i,j}=0\right],$ (8)
here, $z_{i}=1$ indicates an all-zero group (no learning signal) and $z_{i}=0$ indicates at least one non-zero reward (GRPO has exploitable signal). To reduce step-to-step noise, we maintain an exponential moving average (EMA):
$\tilde{z}_{i}^{(t)}\;=\;\alpha\,\tilde{z}_{i}^{(t-1)}\;+\;(1-\alpha)\,z_{i}^{(t)},\qquad\alpha\in[0,1),$ (9)
computed per prompt (or per minibatch) at step $t$.
We then map $\tilde{z}_{i}^{(t)}$ to a bounded, monotonic mixing weight that will later control how much oracle education to apply:
$w_{i}^{(t)}=\mathrm{clip}\Big(w_{\min}+(w_{\max}-w_{\min})\big[\tilde{z}_{i}^{(t)}\big]^{\gamma},w_{\min},w_{\max}\Big),$ (10)
where $w_{\min}\!\leq\!w_{i}^{(t)}\!\leq\!w_{\max}$ enforces a narrow operating range for stability. The exponent $\gamma$ shapes sensitivity: larger $\gamma$ concentrates education on the most severe failures while leaving ordinary prompts to GRPO.
Integrating Two Objectives. With the ZRR-derived weight $w_{i}^{(t)}$ in hand, we modulate learning per prompt: when GRPO provides usable signal ($w_{i}^{(t)}$ is small), we emphasise exploration and exploitation; when GRPO stalls ($w_{i}^{(t)}$ is large), we shift toward direct education via preference alignment. We combine both signals in a single objective over a batch of size B:
$\mathcal{L}_{\mathrm{OraPO}}=\frac{1}{B}\sum_{i=1}^{B}\Big[\big(1-w_{i}^{(t)}\big)$ $\mathcal{L}_{\mathrm{GRPO}}(x_{i},p_{i})$ (11)
$+w_{i}^{(t)}$ $\mathcal{L}_{\mathrm{DPO}}(x_{i},p_{i})\Big],$
where $\mathcal{L}_{\mathrm{GRPO}}(x_{i},p_{i})$ is the GRPO loss in Eq. (4) evaluated on the sampled completions $(x_{i},p_{i})$, and $\mathcal{L}_{\mathrm{DPO}}(x_{i},p_{i})$ is the DPO loss in Eq. (7) computed on preference pairs associated with the same prompt $(x_{i},p_{i})$. We set the positive to the ground-truth report $y_{i}^{\star}$ and take negatives from all other model completions:
$y_{i}^{+}=y_{i}^{\star},\qquad y_{i}^{-}\in\{\hat{y}_{i,j}\}_{j=1}^{K}\sim\pi_{\theta}(\cdot\mid x_{i},p_{i})$ (12)
where $\{\hat{y}_{i,j}\}_{j=1}^{K}$ is the entire group sampled from $\pi_{\theta}(\cdot\mid x_{i},p_{i})$. We then instantiate $\mathcal{L}_{\mathrm{DPO}}(x_{i},p_{i})$ by applying Eq. (7) over pairs $\{(x_{i},p_{i},y_{i}^{+},y_{i}^{-})\}$. This construction reuses low-reward GRPO generation as a preference negative rather than discarding them. As a result, it enables DPO training without requiring any additional hard-negative mining or auxiliary generators. Generated rollouts with zero or low rewards serve as high-quality, domain-specific negative samples that are otherwise difficult to obtain manually.
Intuitively, a high ZRR flags failure cases where preference education must lead. OraPO responds by increasing $w_{i}^{(t)}$ and applying DPO updates that teach the policy to reject the sampled negatives and align with the ground-truth report. This converts groups into labeled preferences and supplies gradients exactly when GRPO has none/few. A low ZRR indicates that GRPO is uncovering usable reward. This can happen on simpler studies, and it also emerges after DPO has taught the model the cues for difficult findings (e.g., small apical pneumothorax or subtle interstitial edema), enabling a shift from education to exploration and exploitation. In practice, we observe the intended training dynamic, as shown in Fig. 2: early training is education-heavy with larger ratio of zero-reward groups $\rightarrow w^{(t)}$, then $w^{(t)}$ gradually declines as the policy learns and GRPO takes over.
What does OraPO Buy Us? As shown in Fig. 2, OraPO drives down the zero-reward ratio faster, and to a lower plateau than naïve GRPO, implying fewer wasted rollouts and more gradient-carrying updates. Corroborating this, the class-level performance curves for Pneumonia and Fracture (clinically challenging, low-prevalence labels) show that OraPO improves earlier and sustains higher F1 across checkpoints. The coupled trend (rapid ZRR decline $\rightarrow$ earlier F1 lift on hard classes) indicates that DPO is converting failed exploration into effective supervision rather than merely stabilising optimisation, yielding both better sample efficiency and stronger performance where it matters.
This initiates a self-reinforcing data flywheel: as the model learns through Oracle education and GRPO’s exploration–exploitation process, it produces increasingly informative negative samples. These, in turn, carry higher reward signals, pushing the model to achieve even greater rewards and completing a positive, compounding feedback cycle.
3.3 FactScore-based Reward Design
OraPO needs a reward that is reliable and efficient to compute. Many RL-based RRG approaches optimize report-level proxies or embedding similarity, which favors fluent paraphrases over verifiable clinical facts and can yield plausible but clinically incorrect narratives. We introduce a FactScore-based reward (FactS) that treats the report as its own rationale: extract atomic clinical facts and test each against the ground-truth label set to produce a dense, fact-entailment signal for reinforcement learning.
Step 1: Extract atomic clinical facts. Inspired by FactScore [42], we define a fact as an atomic, verifiable clinical statement (e.g. finding, attribute, and optionally location), for instance, “no pleural effusion”, “linear atelectasis at the left base”. Given a generated report $\hat{y}_{i}$, we prompt GPT-4.1 [49] to extract a set of atomic facts:
$\mathcal{F}(\hat{y}_{i})=\{s_{i,k}\}_{k=1}^{K_{i}},$ (13)
ensuring each $s_{i,k}$ is concise and verifiable.
Step 2: Per-label entailment against ground-truth. Let $\mathcal{L}$ be the fixed label set (e.g. 14 CheXpert labels) and let $z^{\ast}_{i,\ell}\in\{0,1\}$ be the ground-truth for label $\ell\in\mathcal{L}$ on study $x_{i}$. We then run an LLM-based entailment check that maps the fact set to label-level predictions:
$\hat{z}_{i,\ell}\;=\;\mathbf{1}\!\left[\exists\,s\in\mathcal{F}(\hat{y}_{i})\ \text{s.t.}\ \texttt{entails}(s,\ell)=\texttt{true}\right]$ (14)
for all $\ell\in\mathcal{L}$. Explicit contradictions to $\ell$ are treated as false positives for that label (discouraging unsupported claims).
Step 3: A dense, interpretable reward. Based on $(\hat{z}_{i,\ell},z^{\ast}_{i,\ell})_{\ell\in\mathcal{L}}$, we compute standard per-instance precision $P_{i}$ and recall $R_{i}$, and the reward as an $F_{\beta}$ score
$r(x_{i},\hat{y}_{i})=F_{\beta,i}=\frac{(1+\beta^{2})\,P_{i}R_{i}}{\beta^{2}P_{i}+R_{i}+\xi},$ (15)
with $\beta>1$ placing more weight on recall (penalizing missed positive findings) and $\beta<1$ emphasizing precision; $\xi$ is a small constant for stability. The reward is dense and label-wise interpretable, directly supervising diagnostic correctness while remaining token-efficient. It requires no critic, no auxiliary reward model, nor long reasoning traces. By anchoring each sentence to specific clinical facts and aligning them with ground-truth labels, the FactS Reward reduces unsupported claims and improves coverage of required findings, providing a faithful signal for OraPO.
Table 1: Experimental results on the CheXpert Plus dataset [8] using RRG task-specific algorithms. We report per-label Precision, Recall, and F1 (14 pathologies). Train Size is the number of training samples. Some baselines train on multiple corpora and/or in multiple stages. Best and second best are in bold and underline, respectively.
Algorithm Venue Precision, Recall, F1 Train Size
R2GenRL [51] ACL22 0.193, 0.229, 0.196 223K
XProNet [57] ECCV22 0.314, 0.247, 0.259 223K
MSAT [63] MICCAI22 0.044, 0.142, 0.057 223K
ORGan [25] ACL23 0.288, 0.287, 0.277 223K
M2KT [70] MIA21 0.044, 0.142, 0.058 223K
TIMER [66] CHIL23 0.345, 0.238, 0.234 223K
CvT2DistilGPT2 [46] AIM23 0.285, 0.252, 0.246 223K
R2Gen [13] EMNLP20 0.318, 0.200, 0.181 223K
R2GenCMN [14] ACL21 0.329, 0.241, 0.231 223K
Zhu et al. [81] MICCAI23 0.217, 0.308, 0.205 223K
CAMANet [58] IEEE JBHI23 0.328, 0.224, 0.216 223K
ASGMD [68] ESWA24 0.146, 0.108, 0.055 223K
Token-Mixer [71] IEEE TMI23 0.309, 0.270, 0.288 223K
PromptMRG [27] AAAI24 0.258, 0.265, 0.281 223K
R2GenGPT [65] Meta-Rad.23 0.315, 0.244, 0.260 223K
WCL [69] EMNLP21 0.335, 0.259, 0.256 223K
R2GenCSR [60] arXiv24 0.315, 0.247, 0.259 223K
Wang et al. [61] VIS INT25 0.175, 0.099, 0.078 223K
CheXagent [15] arXiv24 0.373, 0.194, 0.186 8.5M
MambaXray-B [62] CVPR25 0.333, 0.264, 0.273 1.27M
MambaXray-L [62] CVPR25 0.377, 0.319, 0.335 1.27M
VLCI [12] IEEE TIP25 0.341, 0.175, 0.163 223K
Ours 0.237, 0.832, 0.341 1K
Table 2: Experimental Results on the CheXpert Plus benchmark [8] across diverse LLMs/VLMs, all supervised fine-tuned with R2GenGPT [65] on CheXpert Plus. We report macro Precision, Recall, and F1 (14 pathologies). The Params listed denotes the parameters that need to be tuned in the training phase. FT-Size indicates the number of samples used for fine-tuning. Best and second best are in bold and underline, respectively.
LLM/VLM Year Precision, Recall, F1 Params FT-Size
Vicuna-V1.5 [78] 2023 0.334, 0.258, 0.276 6.7B 223K
Qwen-1.5 [3] 2024 0.303, 0.233, 0.241 7.7B 223K
Qwen-2 [3] 2024 0.313, 0.269, 0.261 7.6B 223K
InternLM [7] 2024 0.307, 0.274, 0.284 7.3B 223K
Llama-2 (7B) [56] 2023 0.315, 0.244, 0.260 6.7B 223K
Llama-2 (13B) [56] 2023 0.321, 0.254, 0.267 13.0B 223K
Llama-3 [18] 2024 0.306, 0.232, 0.222 8.0B 223K
Llama-3.1 [18] 2024 0.295, 0.251, 0.242 8.0B 223K
GPT2-Medium [52] 2019 0.358, 0.186, 0.165 354M 223K
Orca-2 (7B) [43] 2023 0.330, 0.251, 0.271 6.7B 223K
Orca-2 (13B) [43] 2023 0.317, 0.242, 0.257 13.0B 223K
Deepseek-LLM [5] 2024 0.336, 0.256, 0.253 6.9B 223K
Yi-1.5 (6B) [73] 2024 0.322, 0.229, 0.226 6.1B 223K
Yi-1.5 (9B) [73] 2024 0.336, 0.241, 0.243 8.8B 223K
InternVL-2 [7] 2023 0.196, 0.127, 0.132 8.0B 223K
MiniCPM-V2.5 [72] 2024 0.254, 0.152, 0.122 8.4B 223K
Ours 0.237, 0.832, 0.341 3B 1K
Table 3: Ablation on CheXpert Plus [8]: Impact of the FactS reward and OraPO. The first row indicates the base Qwen2.5-VL-3B-Instruct [4] direct test results, without any fine-tuning. The second row corresponds to a naïve GRPO-based RLHF baseline using accuracy as the reward [21]. We report per-label Precision, Recall, and F1 across 14 findings.
FactS OraPO Train Size Precision, Recall, F1
$\times$ $\times$ 0 0.097,  0.104,  0.034
$\times$ $\times$ 1, 000 0.026,  0.162,  0.089
$\checkmark$ $\times$ 1, 000 0.204,  0.605,  0.291
$\checkmark$ $\checkmark$ 400 0.217,  0.732,  0.296
$\checkmark$ $\checkmark$ 1, 000 0.237,  0.832,  0.341
4 Experiments
4.1 Experimental Setup
Datasets. CheXpert Plus [8] is a recent large-scale paired image–text resource that links chest X-ray DICOMs to fully de-identified radiology reports, organized into report subsections and annotated for 14 target pathologies (i.e. disease labels). The public release includes $\sim$223K image–report pairs across $\sim$187.7k studies and $\sim$64.7k patients, totaling $\sim$36M report tokens. For data-efficient reinforcement fine-tuning, we subsample a pool of 1,000 studies with a balanced ground-truth label set via CheXpert labeller [26]. We also exclude reports with missing sections. Following recent works [62], we report results on the test split and extract labels from both generated and GT reports by the CheXpert labeller.
Baselines. Following the recent CheXpert Plus benchmark study [62], we compare our proposed approach against a broad suite of methods: 22 task-specific RRG models spanning diverse architectures: ORGan [25], M2KT [70], ASGMD [68], Token-Mixer [71], PromptMRG [27], R2GenRL [51], XProNet [57], MSAT [63], TIMER [66], CvT2DistilGPT2 [46], R2Gen [13], R2GenCMN [14], Zhu et al. [81], CAMANet [58], R2GenGPT [65], WCL [69], VLCI [12], Wang et al. [61], R2GenCSR [60], and the recent MambaXray-B/L [62], as well as CheXagent [15], and 16 open-source LLMs/VLMs, including Vicuna-7B [78], Qwen1.5-7B and Qwen2-7B-Instruct [3], InternLM-7B [7], Llama-2 (7B/13B) [56], Llama-3/3.1 (8B) [18], GPT-2 Medium [52], Orca-2 (7B/13B) [43], DeepSeek-LLM-7B-Chat [5], Yi-1.5 (6B/9B) [73], InternVL-2 [7], and MiniCPM-V2.5 [72]. All task-specific algorithms are fine-tuned on at least the full CheXpert Plus training set, and some of them on multi-source corpora. Thus, our 1K training set represents $\sim$0.45% of the standard 223K split; and only $\sim$0.079% of the 1.27M used by MambaXray [12], which is the current SOTA model.
Evaluation Metric. We follow [62] to report clinical effectiveness as macro-averaged over labels: for each of the 14 conditions we compute Precision, Recall, and F1 on the test set, then take the unweighted mean across labels so frequent and rare conditions contribute equally.
Implementation Details.
We fine-tune a 3B Qwen-2.5-VL policy [4] on the 1,000 CheXpert Plus studies in an RL-only setting (no SFT/alignment), using Hugging Face TRL’s GRPOTrainer with transformers and accelerate. Training runs on four NVIDIA A10 GPUs (24 GB) with PyTorch 2.7.1/CUDA 12.1. We empirically set batch size $B=16$. For each prompt, we sample $K=8$ completions, compute the FactS reward per completion, and optimize the objective in Eq. (11): GRPO provides the default learning signal, while DPO is invoked adaptively via the ZRR schedule. To prioritize GRPO yet correct hard, zero-reward cases, we map the EMA zero-reward rate $\tilde{z}_{i}^{(t)}$ with $\alpha=0.5$ in Eq. (9) to a bounded mixing weight $w_{i}^{(t)}$ in Eq. (10) with $w_{\min}\!=\!0.05,\quad w_{\max}\!=\!0.15,\quad\gamma\!=\!2.$ Thus, $w_{i}^{(t)}\!\to\!w_{\max}$ only when a prompt’s ZRR approaches 1 (no usable GRPO signal), and it quickly decays toward $w_{\min}$ once GRPO yields non-zero rewards, ensuring GRPO dominates ordinary updates and DPO remains a light, targeted aid. We adopt two length-aware variants to handle variable-length reports: DR.GRPO [38] to mitigate length bias in on-policy updates, and LN-DPO [2] to normalize preference margins by sequence length.
4.2 Main Results and Analysis on CheXpert Plus
Against RRG task-specific algorithms.
As shown in Tab. 1, OraPO+FactS sets the new SOTA F1 = 0.341 (best) and Recall = 0.832 (best) on CheXpert Plus, significantly surpassing best-performing baselines by $\mathbf{+160.8\%}$ in recall. When compared with the most recent baseline, VLCI, our proposed method outperforms VLCI by $\mathbf{+109.2\%}$ in F1 and $\mathbf{+375.4\%}$ in recall. While our precision trails MambaXray-L, this is a deliberate trade-off between emphasizing per-finding coverage and accurate statements of detected findings. In practice, recall-oriented drafts are preferred, as radiologists perform the final review, they often prefer high-sensitivity assistance [23]. Overall, our method sets the new SOTA performance in F1 and recall, with three orders of magnitude less training data.
Against Alignment-based SFT for LLM/VLM. In this experiment, we compare our framework against R2GenGPT [65], a popular alignment-based SFT approach that is widely used for training large LLM/VLM. In this approach, a visual encoder is aligned with a language model into the same space via supervised fine-tuning on paired image-report data. We report the performance of R2GenGPT [65] with different LLM/VLM backbones in Tab. 2. Our approach is trained with a compact 3B policy on 1K CheXpert Plus samples. Under this small-data setup, OraPO attains $\mathbf{F1=0.341}$ and $\mathbf{Recall=0.832}$, significantly outperforming larger aligned backbones trained on 223K: F1 improves over InternLM by $\mathbf{+20.1\%}$ and over Qwen-2 (0.261) by $\mathbf{+30.7\%}$; recall improves over InternLM by $\mathbf{+203.6\%}$ and over GPT-2 Medium by $\mathbf{+347.3\%}$. These results validate that our proposed OraPO with fact-level supervision is a superior learning paradigm than standard vision–language alignment, achieving higher performance with less than half backbone parameters and markedly less training data.
4.3 Ablation Study
To study the effectiveness and efficiency of our proposed method, we evaluate four model variants on CheXpert Plus using macro Precision, Recall, and F1. As reported in Tab. 3, the second row is a naïve GRPO-based RLHF baseline that uses accuracy as reward (as in DeepSeek-R1 [21]), which yields only marginal gains over the base model without any domain-specific training (row 1). Adding FactS to GRPO (row 3) lifts F1 from 0.089 to 0.291 $(+227\%)$, Recall from 0.162 to 0.605 (+274%), and Precision from 0.026 to 0.204 (+685%), showing that fact-level entailment provides dense, clinically aligned supervision. Layering OraPO on top (row 5) further improves F1 to 0.341 (+17.2% vs. FactS-only), with Recall improved by (+37.5%) and Precision improved by (+16.2%). Notably, OraPO+FactS with only 400 studies (row 4) attains F1 == 0.296, Precision == 0.217 and Recall == 0.732—–surpassing the FactS-only 1K setting (F1 == 0.291, Precision == 0.204, and Recall == 0.605)–—providing evidence of OraPO’s superb data efficiency. Overall, FactS supplies faithful, sentence-level rewards, while OraPO converts zero-reward groups into preference updates, yielding higher precision & recall and earlier learning on scarce yet clinically important signals.
5 Conclusion
In this work, we present OraPO, an oracle-educated RL training algorithm for high-quality radiology report generation under tight data and compute budgets. To our knowledge, we are the first to couple DPO with GRPO: when GRPO exploration fails, we recycle those rollouts as negatives and apply a lightweight DPO step to recover useful gradients. In parallel, we design the FactS reward to provide dense, fact-level, clinically faithful supervision tailored to RRG. Comprehensive experiments show OraPO setting the new SOTA performance while using 2–3 orders of magnitude fewer training samples.
We have identified a number of promising future work directions. (i) While designed for the RRG task, OraPO can be applied to other hard, general-domain tasks where vanilla GRPO struggles with sparse success signals (e.g., multi-step mathematical reasoning and code synthesis). We will extend experiments to other tasks to validate the effectiveness of OraPO. (ii) Due to compute constraints, our results currently are limited to a 3B backbone VLM with only 1K studies for training. We plan to study OraPO’s scaling behavior across model sizes and data/training budgets. (iii) While our reward design already bridges the gap between verifiable and non-verifiable model outputs by leveraging LLMs as a judge, it still falls short of handling genuinely non-verifiable rewards. Inspired by the recent Rubrics as Rewards technique [20], we plan to extend our methodology to address the challenge of non-verifiable reward modeling. (iv) FactS mainly scores disease presence/absence and underweights attributes (location, laterality, size/severity, temporal change). We will design a more fine-grained fact-to-attribute entailment reward for addressing the misalignment issue on diagnostic attributes.