### v2

# OraPO: Oracle-educated Reinforcement Learning for Data-efficient and Factual Radiology Report Generation

Abstract
Radiology report generation (RRG) aims to automatically produce clinically faithful reports from chest X-ray images. Prevailing work typically follows a scale-driven paradigm, by multi-stage training over large paired corpora and oversized backbones, making pipelines highly data- and compute-intensive. In this paper, we propose Oracle-educated GRPO (OraPO) with a FactScore-based reward (FactS) to tackle the RRG task under constrained budgets. OraPO enables single-stage, RL-only training by converting failed GRPO explorations on rare or difficult studies into direct preference supervision via a lightweight oracle step. FactS grounds learning in diagnostic evidence by extracting atomic clinical facts and checking entailment against ground-truth labels, yielding dense, interpretable sentence-level rewards. Together, OraPO and FactS create a compact and powerful framework that significantly improves learning efficiency on clinically challenging cases, setting the new SOTA performance on the CheXpert Plus dataset (0.341 in F1) and MIMIC-CXR dataset (0.357 in F1) with 2–3 orders of magnitude less training data using a small base vision-language model on modest hardware.
Figure 1: Illustration of mainstream data/compute-intensive pipelines (upper-left) versus our data-efficient pipeline (upper-right). Bottom: on CheXpert Plus [8], our method achieves the SOTA performance for RRG with less than 0.1% of the training samples (vs. 1.27 M) used by best-performing baselines and a much smaller model, demonstrating strong performance under tight data and compute budgets.
1 Introduction
Radiology report generation (RRG) from chest X-rays aims to turn imaging evidence into a clinically usable narrative. This is a complex multimodal task that couples disease prediction with free-text report generation, linking detected visual findings (e.g. edema, pleural effusion, and pneumothorax) to clear sentences that describe their presence, details, location, and severity. Automating this task matters in practice: the large imaging volumes and backlogs place serious strains on healthcare services and delay patient care, while radiologist staffing remains insufficient. For example, England reports a 29% radiology consultant shortfall alongside sustained vacancies [68], while in the United States there are about 14,000 openings but only 1,150 resident graduates annually [77]. In this context, AI-assisted drafting has begun to show measurable efficiency gains. A recent study [1] reported that using draft reports generated by large language models (LLMs) reduced reporting time by 24%. Given its practical importance, automated radiology report generation has rapidly become an active research area aimed at easing radiologists’ workload and streamlining workflows in hospitals and clinics.
Mainstream RRG methods typically rely on supervised fine-tuning (SFT) via multi-stage training on large corpora and/or scaling up backbone size (Fig. OraPO: Oracle-educated Reinforcement Learning for Data-efficient and Factual Radiology Report Generation, upper-left). The multi-stage training paradigm usually follows three phases: domain-specific pre-training, image–text alignment, and task-specific fine-tuning, where the first two are usually trained on combinations of large-scale, multi-source datasets such as medical textbooks, real radiology reports including but not limited to MIMIC-CXR [41], IU X-ray [21] and CheXpert Plus [8], often for dozens of epochs [92, 93]. Notably, both CheXpert Plus and MIMIC-CXR include training splits exceeding 200K image-report pairs. In addition, there is a recent trend of adopting large vision-language models (VLMs) with >>13B parameters for SFT, demanding meticulous data curation and substantial GPU budgets [45, 42].
Recently, Group Relative Policy Optimisation (GRPO) has emerged as a strong and compute-efficient RL-based learning algorithm for many hard problems with verifiable rewards. Its efficiency stems from two fronts: (i) lower memory/compute by dispensing with the trainable value critic and instead computing group-normalised advantages from the relative rewards of multiple samples [63, 44]; and (ii) a simplified training recipe that enables “RL-only” regimes (i.e., without costly SFT/alignment), avoiding massive pre-training/alignment corpora [79, 114, 20, 10].
However, applying GRPO to the RRG task reveals two challenges. Challenge (i): we observe that vanilla GRPO often produces all-zero reward groups early in training (e.g. 30% in the first 50 steps, as shown in Fig. 2 left), causing vanishing gradients and wasted rollouts. Recent variants attempt to fix this by resampling until a non-zero reward appears [110] or by enlarging the group size [81], both of which are still suboptimal due to increased cost of more rollouts. Another line interleaves SFT with RL to maintain gradient flow, yet still leaves low-quality rollouts unused, wasting costly sampling [58, 11, 112, 57, 54]. Challenge (ii): vanilla GRPO requires readily verifiable scalar rewards. Unlike in multiple-choice QA, maths, or coding, where the reward is typically a binary final-answer check, radiology reports are long-form, multi-fact, and must be consistent across sentences, making reward design both critical and difficult. In practice, RLHF-style RRG work often uses reference-overlap metrics that mainly capture fluency/surface similarity (e.g. BLEU/CIDEr [48, 67]) or a report-level clinical metric [120, 62, 107]. These proxies under-penalise sentence-level factual errors and cross-sentence contradictions, yielding fluent yet clinically misleading narratives (missed positives, unsupported claims). Hence, a reward must be able to assess factual correctness of each statement of a radiology report in an efficient manner.
In this paper, we propose Oracle-educated Group Relative Policy Optimisation: OraPO, a novel RL-based learning algorithm that tackles RRG’s challenges in a data- and compute-efficient manner. OraPO addresses Challenge (i) of GRPO by dynamically injecting oracle supervision (i.e. ground-truth) for groups with all-zero rewards and teaching the policy to avoid repeating low-quality generations. Specifically, When a sampled group yields all-zero rewards (group 1 in Fig. OraPO: Oracle-educated Reinforcement Learning for Data-efficient and Factual Radiology Report Generation, upper-right), OraPO triggers a lightweight Direct Preference Optimisation (DPO) update that directly prefers the ground-truth report over the zero-reward rollouts, effectively reusing those failed rollouts as ready-made negative examples, with no extra compute or annotation needed. This converts failed exploration in difficult cases (hard or low-prevalence studies) into useful gradients, moving the policy away from low-quality modes and toward the ground-truth. We also introduce a simple adaptive mixing weight to increase DPO’s influence when zero-reward groups are frequent, which gradually tapers once GRPO provides informative advantages, thereby allowing OraPO to salvage otherwise wasted batches, stabilise training, and accelerate convergence. .
Moreover, to provide OraPO with clinically faithful supervision, we introduce a FactScore-based reward [60] designed to address Challenge (ii). We treat the generated report as its own rationale, extract atomic clinical facts, and test their entailment against the ground-truth label set. Facts that support a label contribute positively; unsupported or contradictory labels incur penalties, yielding a dense, interpretable per-label reward. This stands in sharp contrast to the original GRPO reward, which assigns reward solely to the final answer while treating the reasoning trace largely as a byproduct of generation, without ensuring it is factually grounded or consistently aligned with the answer. Our reward design ensures the generated report (serving in lieu of reasoning trace) not only produces a correct cognitive reasoning trace similar to that of an experienced radiologist but also yields correct diagnostic findings, which is crucial in domains such as healthcare in which safety is of paramount importance.
Extensive experiments are conducted on CheXpert Plus [8] and MIMIC-CXR [8] datasets, where OraPO sets the new SOTA performance for RRG, surpassing the previous best model with F1 score and notably achieving 160.8% improvement in recall. This higher recall is clinically meaningful. In automated RRG tasks, recall is often prioritised over precision, as false negatives (missed abnormalities) can delay or prevent timely treatment, whereas false positives generally result only in additional clinical review [62, 78, 76, 69]. Previous studies [43, 107] emphasise maintaining high recall to avoid critical finding omissions, recognising that such oversights carry more serious clinical consequences than reduced precision. The proposed method achieves this clinical effectiveness while being exceptionally data-efficient, using 2–3 orders of magnitude less training data (i.e. trained on only 1K samples compared to the current best baseline MambaXray-L’s 1.27M samples and other baselines’ minimum of 223K samples). At the same time, OraPO is compute-efficient: it is based on Qwen2.5-VL-3B, a small-sized VLM on modest hardware (4× A10 GPUs).
In summary, our contributions are threefold:
  1. 1.
    We present OraPO, an Oracle (DPO) educated GRPO that turns failed explorations into direct preference supervision significantly improving data- and compute-efficiency. To the best of our knowledge, this is the first work to integrate direct preference learning with GRPO-based RL for efficient training.
  2. 2.
    We design a FactS Reward for RL, performing fact-level entailment against ground-truth labels, yielding dense, interpretable feedback that aligns reports with diagnostic facts while avoiding long reasoning traces.
  3. 3.
    We show that OraPO achieves new SOTA performance for RRG, surpassing the previous best model (MambaXray-L) in clinical effectiveness F1 and recall while using 2–3 orders of magnitude less training data.
Figure 2: Left: Cumulative proportion of zero-reward batches (reward batch mean = 0) vs. training step on CheXpert Plus [92]. OraPO suppresses zero-reward frequency faster than naïve GRPO. Centre/Right: Class-level F1 on the CheXpert Plus validation set [92] across checkpoints for two clinically challenging and rare classes: Pneumonia (2.70%) and Fracture (4.05%). OraPO learns earlier and maintains higher F1 than naïve GRPO.
2 Related Work
Automatic Radiology Report Generation (RRG) aims to turn imaging studies (e.g., chest X-rays) into clinically usable free-text reports. Early approaches adopted seq2seq networks (e.g., CNN–RNN and Transformer) [40, 14, 49, 109, 108] for training on large paired corpora such as MIMIC-CXR [41]. Recent work advances three fronts: (i) knowledge- and rule-guided generation, which injects medical priors to curb hallucination [6] and performs disease-aware image–text alignment for trustworthy text [72, 38]; (ii) leveraging prior studies and multi-view images to improve coverage and consistency [30, 53, 17]; and (iii) scaling/standardizing with pre-training, releasing multi-stage pre-training recipes [92, 30, 16, 18, 37]. In parallel, LLM-driven RRG explores instruction-tuned or preference-aligned training [52, 85, 45]. Despite progress, these approaches often rely on large paired corpora and/or costly computation resources, leading to deficiencies in data efficiency and practical scalability.
Group Relative Policy Optimisation (GRPO) is a critic-free, reinforcement learning objective for generative LLMs, introduced in DeepSeekMath [79]. It samples multiple completions per prompt, computes group-normalised advantages from relative rewards, and updates the policy with clipped ratios to a frozen reference; removing the critic reduces memory/compute versus PPO while preserving stability [79, 29, 113, 63, 44]. Recent variants mainly target optimisation stability: DR.GRPO [56] mitigates length bias and skewed group rewards; CISPO [9] clips importance-sampling weights; GSPO [115] adopts sequence-level clipping for MoE; DAPO [110] resamples until a non-zero reward appears. However, these methods were not designed to improve data efficiency: most target optimisation stability while leaving data volume and training epochs essentially unchanged, and some of them often increase sampling/compute [110, 89, 19, 81].
Direct Preference Optimisation (DPO) trains a policy from preference pairs to prefer the chosen response over the rejected one [119]. Recent variants [111] streamline or improve the robustness of this recipe: SimPO [59] applies length-normalised mean log-probabilities with a simplified reference; Cal-DPO [98] calibrates the implicit reward to human scales; ORPO [33] drops the reference via an odds-ratio objective integrated into SFT; and KTO [25] learns from unary thumbs-up/down signals instead of pairs. In our RRG framework, DPO naturally serves as an oracle step inside OraPO to convert those failed GRPO rollouts into useful preference updates against the ground-truth, leading to data- and compute-efficient learning.
3 Methodology
In this section, we first introduce the background of Group Relative Policy Optimisation (GRPO) and Direct Preference Optimisation (DPO) (§3.1). We then identify a failure mode of GRPO in radiology report generation (RRG), when a pretrained model produces many low-quality groups (e.g., zero rewards) that yield little or no gradient. This observation motivates our proposed OraPO algorithm, which efficiently converts such failed explorations into preference supervision (§3.2). Finally, we design the FactS Reward, a fact-level entailment signal that aligns reports with diagnostic labels and provides dense, interpretable feedback for reinforcement learning (§3.3).
3.1 Preliminaries
Group Relative Policy Optimisation (GRPO). GRPO [79, 29] is a computation- and memory-efficient RL objective for aligning generative models (LLMs/VLMs), introduced for reasoning-heavy settings and now used broadly because of its efficient critic-free design and proven effectiveness. It replaces the value critic with group-normalized advantages computed among multiple completions sampled for the same input. Given a study–prompt pair $(x,p)$, the policy $\pi_{\theta}$ draws a group of $K$ sequences $\{\hat{y}_{j}\}_{j=1}^{K}\sim\pi_{\theta}(\cdot\mid x,p)$ and receives scalar rewards $r_{j}=r(x,\hat{y}_{j})$. GRPO forms a group baseline and scale,
$\bar{r}=\frac{1}{K}\sum_{j=1}^{K}r_{j},\qquad\sigma=\sqrt{\frac{1}{K}\sum_{j=1}^{K}(r_{j}-\bar{r})^{2}+\varepsilon},$ (1)
and defines the group-normalised advantage:
$A_{j}=\frac{r_{j}-\bar{r}}{\sigma}.$ (2)
Let the proximal policy optimisation (PPO) ratio be:
$\rho_{j}=\exp\!\big(\log\pi_{\theta}(\hat{y}_{j}\mid x,p)-\log\pi_{\theta_{\text{old}}}(\hat{y}_{j}\mid x,p)\big),$ (3)
where $\theta_{\text{old}}$ denotes the parameters of the behaviour policy used to sample the current group (i.e., the parameters before the update). Then the GRPO loss over a batch is defined as:
$\mathcal{L}_{\mathrm{GRPO}}$ $=-\,\mathbb{E}\Big[\min\big(\rho_{j}A_{j},\ \mathrm{clip}(\rho_{j},1-\epsilon,1+\epsilon)\,A_{j}\big)\Big]$ (4)
$\quad+\lambda_{\mathrm{KL}}\ \mathbb{E}\!\left[\mathrm{KL}\!\big(\pi_{\theta}(\cdot\mid x,p)\,\|\,\pi_{\mathrm{ref}}(\cdot\mid x,p)\big)\right],$
where $\epsilon$ is the clipping threshold, $\lambda_{\mathrm{KL}}\geq 0$ controls regularization toward a frozen reference policy $\pi_{\mathrm{ref}}$. In practice, GRPO samples $K$ completions per prompt, computes $r_{j}$, normalizes within-group to obtain $A_{j}$, and updates $\pi_{\theta}$ with the clipped-ratio objective in (4), thereby achieving critic-free policy optimisation.
Direct Preference Optimisation (DPO). DPO [119] is a widely used, preference-alignment objective for generative LLMs/VLMs. It updates the policy to prefer chosen completions over rejected ones relative to a frozen reference model. Given a preference dataset $\mathcal{P}=\{(x,p,y^{+},y^{-})\}$, where $y^{+}$ is preferred to $y^{-}$ for the same input–prompt pair $(x,p)$, DPO maximizes the log-probability margin of the preferred sample after subtracting the reference log-probabilities. Let $\pi_{\theta}$ be the trainable policy and $\pi_{\mathrm{ref}}$ the frozen reference. The expectation below is taken over preference pairs drawn from $\mathcal{P}$. With temperature $\tau>0$, define
$\Delta^{+}=\log\pi_{\theta}(y^{+}\mid x,p)\;-\;\log\pi_{\mathrm{ref}}(y^{+}\mid x,p),$ (5)
$\Delta^{-}=\log\pi_{\theta}(y^{-}\mid x,p)\;-\;\log\pi_{\mathrm{ref}}(y^{-}\mid x,p),$ (6)
and the DPO loss
$\mathcal{L}_{\mathrm{DPO}}=-\,\mathbb{E}\Big[\log\sigma\!\big(\tau\,(\Delta^{+}-\Delta^{-})\big)\Big],$ (7)
where $\sigma(\cdot)$ is the logistic sigmoid.
3.2 OraPO: Oracle-educated GRPO
Exploration Failure of GRPO in RRG. On radiology report generation, applying vanilla GRPO to a base VLM with limited radiology knowledge often fails to explore effectively. The root reason is that the base model carries very limited domain knowledge about radiology. Most base VLMs [117] are pre-trained on general vision–language corpora and tasks, for example, everyday-image captioning on MS COCO [50, 96] and open-domain visual question answering on VQA v2 [26], rather than medical imaging and reporting. Hence, it tends to miss or misstate rare and clinically challenging findings (e.g., small pneumothorax, subtle interstitial edema, or linear atelectasis) when asked to produce clinical narratives. For many chest X-ray inputs, the sampled group of completions is uniformly low quality with near-zero rewards. Thus, the policy receives little learning signal to drive effective exploration.
Empirically, before any domain-specific fine-tuning, the pretrained backbone attains only $0.034$ macro-F1 (Tab. 4) on clinical effectiveness. As shown in Fig. 2 (left), in the first $50$ optimisation steps approximately $30\%$ of groups have all rewards equal to zero. Under the GRPO construction, such groups give $\bar{r}\approx 0$ and $\sigma\approx\varepsilon$, so $A_{j}=\tfrac{r_{j}-\bar{r}}{\sigma}\approx 0$ for all $j$. Consequently, the clipped-ratio term in Eq. (4) contributes essentially no gradient, and the update is dominated by the KL regularizer to $\pi_{\mathrm{ref}}$ rather than by task reward. In effect, the optimizer wastes GPU resources by sampling $K$ completions per prompt without extracting useful signals, which both slows convergence and squanders valuable compute.
Zero-Reward Rate (ZRR). To resolve GRPO’s exploration failure on RRG, we first detect when a prompt yields little or no usable signal, then adapt learning accordingly. We introduce a Zero-Reward Rate (ZRR) that quantifies, per prompt, how severely GRPO has stalled.
For prompt $x_{i}$ with a GRPO group of $K$ completions $\{\hat{y}_{i,j}\}_{j=1}^{K}$ and rewards $r_{i,j}\in[0,1]$, define the raw zero-reward rate:
$z_{i}\;=\;\frac{1}{K}\sum_{j=1}^{K}\mathbf{1}\!\left[r_{i,j}=0\right],$ (8)
here, $z_{i}=1$ indicates an all-zero group (no learning signal) and $z_{i}=0$ indicates at least one non-zero reward (GRPO has exploitable signal). To reduce step-to-step noise, we maintain an exponential moving average (EMA):
$\tilde{z}_{i}^{(t)}\;=\;\alpha\,\tilde{z}_{i}^{(t-1)}\;+\;(1-\alpha)\,z_{i}^{(t)},\qquad\alpha\in[0,1),$ (9)
computed per prompt (or per minibatch) at step $t$.
We then map $\tilde{z}_{i}^{(t)}$ to a bounded, monotonic mixing weight that will later control how much oracle education to apply:
$w_{i}^{(t)}=\mathrm{clip}\Big(w_{\min}+(w_{\max}-w_{\min})\big[\tilde{z}_{i}^{(t)}\big]^{\gamma},w_{\min},w_{\max}\Big),$ (10)
where $w_{\min}\!\leq\!w_{i}^{(t)}\!\leq\!w_{\max}$ enforces a narrow operating range for stability. The exponent $\gamma$ shapes sensitivity: larger $\gamma$ concentrates education on the most severe failures while leaving ordinary prompts to GRPO.
Integrating Two Objectives. With the ZRR-derived weight $w_{i}^{(t)}$ in hand, we modulate learning per prompt: when GRPO provides usable signal ($w_{i}^{(t)}$ is small), we emphasise exploration and exploitation; when GRPO stalls ($w_{i}^{(t)}$ is large), we shift toward direct education via preference alignment. We combine both signals in a single objective over a batch of size B:
$\mathcal{L}_{\mathrm{OraPO}}=\frac{1}{B}\sum_{i=1}^{B}\Big[\big(1-w_{i}^{(t)}\big)$ $\mathcal{L}_{\mathrm{GRPO}}(x_{i},p_{i})$ (11)
$+w_{i}^{(t)}$ $\mathcal{L}_{\mathrm{DPO}}(x_{i},p_{i})\Big],$
where $\mathcal{L}_{\mathrm{GRPO}}(x_{i},p_{i})$ is the GRPO loss in Eq. (4) evaluated on the sampled completions $(x_{i},p_{i})$, and $\mathcal{L}_{\mathrm{DPO}}(x_{i},p_{i})$ is the DPO loss in Eq. (7) computed on preference pairs associated with the same prompt $(x_{i},p_{i})$. We set the positive to the ground-truth report $y_{i}^{\star}$ and take negatives from all other model completions:
$y_{i}^{+}=y_{i}^{\star},\qquad y_{i}^{-}\in\{\hat{y}_{i,j}\}_{j=1}^{K}\sim\pi_{\theta}(\cdot\mid x_{i},p_{i})$ (12)
where $\{\hat{y}_{i,j}\}_{j=1}^{K}$ is the entire group sampled from $\pi_{\theta}(\cdot\mid x_{i},p_{i})$. We then instantiate $\mathcal{L}_{\mathrm{DPO}}(x_{i},p_{i})$ by applying Eq. (7) over pairs $\{(x_{i},p_{i},y_{i}^{+},y_{i}^{-})\}$. This construction reuses low-reward GRPO generation as a preference negative rather than discarding them. As a result, it enables DPO training without requiring any additional hard-negative mining or auxiliary generators. Generated rollouts with zero or low rewards serve as high-quality, domain-specific negative samples that are otherwise difficult to obtain manually.
Intuitively, a high ZRR flags failure cases where preference education must lead. OraPO responds by increasing $w_{i}^{(t)}$ and applying DPO updates that teach the policy to reject the sampled negatives and align with the ground-truth report. This converts groups into labeled preferences and supplies gradients exactly when GRPO has none/few. A low ZRR indicates that GRPO is uncovering usable reward. This can happen on simpler studies, and it also emerges after DPO has taught the model the cues for difficult findings (e.g., small apical pneumothorax or subtle interstitial edema), enabling a shift from education to exploration and exploitation. In practice, we observe the intended training dynamic, as shown in Fig. 2: early training is education-heavy with larger ratio of zero-reward groups, higher $w^{(t)}$ is implied, then $w^{(t)}$ gradually declines as the policy learns and GRPO takes over.
What does OraPO Buy Us? As shown in Fig. 2, OraPO drives down the zero-reward ratio faster, and to a lower plateau than naïve GRPO, implying fewer wasted rollouts and more gradient-carrying updates. Corroborating this, the class-level performance curves for Pneumonia and Fracture (clinically challenging, low-prevalence labels) show that OraPO improves earlier and sustains higher F1 across checkpoints. The coupled trend (rapid ZRR decline $\rightarrow$ earlier F1 lift on hard classes) indicates that DPO is converting failed exploration into effective supervision rather than merely stabilising optimisation, yielding both better sample efficiency and stronger performance where it matters.
This initiates a self-reinforcing data flywheel: as the model learns through Oracle education and GRPO’s exploration–exploitation process, it produces increasingly informative negative samples. These, in turn, carry higher reward signals, pushing the model to achieve even greater rewards and completing a positive, compounding feedback cycle.
3.3 FactScore-based Reward Design
OraPO needs a reward that is reliable and efficient to compute. Many RL-based RRG approaches optimize report-level proxies or embedding similarity, which favors fluent paraphrases over verifiable clinical facts and can yield plausible but clinically incorrect narratives. We introduce a FactScore-based reward (FactS) that treats the report as its own rationale: extract atomic clinical facts and test each against the ground-truth label set to produce a dense, fact-entailment signal for reinforcement learning.
Step 1: Extract atomic clinical facts. Inspired by FactScore [60], we define a fact as an atomic, verifiable clinical statement (e.g. finding, attribute, and optionally location), for instance, “no pleural effusion”, “linear atelectasis at the left base”. Given a generated report $\hat{y}_{i}$, we prompt GPT-4.1 [70] to extract a set of atomic facts:
$\mathcal{F}(\hat{y}_{i})=\{s_{i,k}\}_{k=1}^{K_{i}},$ (13)
ensuring each $s_{i,k}$ is concise and verifiable.
Step 2: Per-label entailment against ground-truth. Let $\mathcal{L}$ be the fixed label set (e.g. 14 CheXpert labels) and let $z^{\ast}_{i,\ell}\in\{0,1\}$ be the ground-truth for label $\ell\in\mathcal{L}$ on study $x_{i}$. We then run an LLM-based entailment check that maps the fact set to label-level predictions:
$\hat{z}_{i,\ell}\;=\;\mathbf{1}\!\left[\exists\,s\in\mathcal{F}(\hat{y}_{i})\ \text{s.t.}\ \texttt{entails}(s,\ell)=\texttt{true}\right]$ (14)
for all $\ell\in\mathcal{L}$. Explicit contradictions to $\ell$ are treated as false positives for that label (discouraging unsupported claims).
Step 3: A dense, interpretable reward. Based on $(\hat{z}_{i,\ell},z^{\ast}_{i,\ell})_{\ell\in\mathcal{L}}$, we compute standard per-instance precision $P_{i}$ and recall $R_{i}$, and the reward as an $F_{\beta}$ score
$r(x_{i},\hat{y}_{i})=F_{\beta,i}=\frac{(1+\beta^{2})\,P_{i}R_{i}}{\beta^{2}P_{i}+R_{i}+\xi},$ (15)
with $\beta>1$ placing more weight on recall (penalizing missed positive findings) and $\beta<1$ emphasizing precision; $\xi$ is a small constant for stability. The reward is dense and label-wise interpretable, directly supervising diagnostic correctness while remaining token-efficient. It requires no critic, no auxiliary reward model, nor long reasoning traces. By anchoring each sentence to specific clinical facts and aligning them with ground-truth labels, the FactS Reward reduces unsupported claims and improves coverage of required findings, providing a faithful signal for OraPO.
Table 1: Experimental results on the CheXpert Plus dataset [8] using task-specific RRG algorithms. We report macro-averaged Precision, Recall, and F1 across 14 CheXpert pathologies. Train Size is the number of training samples. Some baselines train on multiple corpora and/or in multiple stages. Best and second best are in bold and underline, respectively.
Algorithm Venue Precision, Recall, F1 Train Size
R2GenRL [74] ACL22 0.193, 0.229, 0.196 223K
XProNet [87] ECCV22 0.314, 0.247, 0.259 223K
MSAT [95] MICCAI22 0.044, 0.142, 0.057 223K
ORGan [34] ACL23 0.288, 0.287, 0.277 223K
M2KT [103] MIA21 0.044, 0.142, 0.058 223K
TIMER [97] CHIL23 0.345, 0.238, 0.234 223K
CvT2DistilGPT2 [66] AIM23 0.285, 0.252, 0.246 223K
R2Gen [14] EMNLP20 0.318, 0.200, 0.181 223K
R2GenCMN [13] ACL21 0.329, 0.241, 0.231 223K
Zhu et al. [121] MICCAI23 0.217, 0.308, 0.205 223K
CAMANet [88] IEEE JBHI23 0.328, 0.224, 0.216 223K
ASGMD [101] ESWA24 0.146, 0.108, 0.055 223K
Token-Mixer [104] IEEE TMI23 0.309, 0.270, 0.288 223K
PromptMRG [39] AAAI24 0.258, 0.265, 0.281 223K
R2GenGPT [94] Meta-Rad.23 0.315, 0.244, 0.260 223K
WCL [102] EMNLP21 0.335, 0.259, 0.256 223K
R2GenCSR [90] arXiv24 0.315, 0.247, 0.259 223K
Wang et al. [91] VIS INT25 0.175, 0.099, 0.078 223K
CheXagent [15] arXiv24 0.373, 0.194, 0.186 8.5M
MambaXray-B [92] CVPR25 0.333, 0.264, 0.273 1.27M
MambaXray-L [92] CVPR25 0.377, 0.319, 0.335 1.27M
VLCI [12] IEEE TIP25 0.341, 0.175, 0.163 223K
Ours 0.237, 0.832, 0.341 1K
Table 2: Experimental results on the CheXpert Plus benchmark [8] across diverse LLMs/VLMs, all supervised fine-tuned with R2GenGPT [94] on CheXpert Plus. We report macro-averaged Precision, Recall, and F1 across 14 CheXpert pathologies. The Params listed denotes the parameters that need to be tuned in the training phase. FT-Size indicates the number of samples used for fine-tuning. Best and second best are in bold and underline, respectively.
LLM/VLM Year Precision, Recall, F1 Params FT-Size
Vicuna-V1.5 [118] 2023 0.334, 0.258, 0.276 6.7B 223K
Qwen-1.5 [3] 2024 0.303, 0.233, 0.241 7.7B 223K
Qwen-2 [3] 2024 0.313, 0.269, 0.261 7.6B 223K
InternLM [7] 2024 0.307, 0.274, 0.284 7.3B 223K
Llama-2 (7B) [84] 2023 0.315, 0.244, 0.260 6.7B 223K
Llama-2 (13B) [84] 2023 0.321, 0.254, 0.267 13.0B 223K
Llama-3 [23] 2024 0.306, 0.232, 0.222 8.0B 223K
Llama-3.1 [23] 2024 0.295, 0.251, 0.242 8.0B 223K
GPT2-Medium [75] 2019 0.358, 0.186, 0.165 354M 223K
Orca-2 (7B) [61] 2023 0.330, 0.251, 0.271 6.7B 223K
Orca-2 (13B) [61] 2023 0.317, 0.242, 0.257 13.0B 223K
Deepseek-LLM [5] 2024 0.336, 0.256, 0.253 6.9B 223K
Yi-1.5 (6B) [106] 2024 0.322, 0.229, 0.226 6.1B 223K
Yi-1.5 (9B) [106] 2024 0.336, 0.241, 0.243 8.8B 223K
InternVL-2 [7] 2023 0.196, 0.127, 0.132 8.0B 223K
MiniCPM-V2.5 [105] 2024 0.254, 0.152, 0.122 8.4B 223K
Ours 0.237, 0.832, 0.341 3B 1K
Table 3: Experimental results on the MIMIC-CXR dataset [8]. We report macro-averaged Precision, Recall, and F1 across 14 CheXpert pathologies. Train Size is the number of training samples. Some baselines train on multiple corpora and/or in multiple stages. Best are in bold.
Algorithm Venue Precision, Recall, F1 Train Size
R2Gen [14] EMNLP20 0.333, 0.273, 0.276 227K
CXR-RePaiR [24] ML4H21 – , – , 0.274 227K
MRG/SILC [51] TIP24 0.457, 0.337, 0.330 227K
CXRMate [65] IMU24 0.277, 0.351, 0.283 227K
HERGen [86] ECCV24 0.415, 0.301, 0.317 145K
ChEX [64] ECCV24 – , – , 0.326 227K
MambaXray-L [92] CVPR25 0.371, 0.321, 0.340 1.27M
GIT-CXR [82] INFO25 – , – , 0.327 227K
MCA-RG [100] MICCAI25 0.443, 0.306, 0.335 227K
Ours 0.242, 0.891, 0.357 1K
Table 4: Ablation on CheXpert Plus [8]: Impact of the FactS reward and OraPO. The first row indicates the base Qwen2.5-VL-3B-Instruct [4] direct test results, without any fine-tuning. The second row corresponds to a naïve GRPO-based RLHF baseline using accuracy as the reward [29]. We report per-label Precision, Recall, and F1 across 14 findings. Best are in bold.
FactS GRPO DPO SFT Train Size Precision, Recall, F1
$\times$ $\times$ $\times$ $\times$ 0 0.097,  0.104,  0.034
$\times$ $\checkmark$ $\times$ $\times$ 1,000 0.026,  0.162,  0.089
$\checkmark$ $\checkmark$ $\times$ $\times$ 1,000 0.204,  0.605,  0.291
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 400 0.217,  0.732,  0.296
$\checkmark$ $\checkmark$ $\times$ $\checkmark$ 1,000 0.171,  0.176,  0.106
$\checkmark$ $\checkmark$ $\checkmark$ $\times$ 1,000 0.237,  0.832,  0.341
Table 5: Experimental results on CheXpert validation set with gold labels from certified radiologists [36].
Method Train Size Precision, Recall, F1
GPT4.1 [70] 0.292, 0.305, 0.253
GPT5 (Thinking) [71] 0.261, 0.421, 0.301
CheXagent [15] 8.5M 0.462, 0.226, 0.254
MambaXray-B [92] 1.27M 0.362, 0.254, 0.238
MambaXray-L [92] 1.27M 0.395, 0.326, 0.280
Ours 1K 0.234, 0.641, 0.288
4 Experiments
4.1 Experimental Setup
Datasets. CheXpert Plus [8] is a recent large-scale paired image–text resource that links chest X-ray DICOMs to fully de-identified radiology reports, organized into report subsections and annotated for 14 target pathologies (i.e. disease labels). The public release includes $\sim$223K image–report pairs across $\sim$187.7K studies and $\sim$64.7k patients, totaling $\sim$36M report tokens. MIMIC-CXR [41] is a large, publicly available paired image–text resource of chest X-rays in DICOM format with contemporaneous, de-identified radiology reports. It contains 377,110 images from 227,943 imaging studies covering 65,379 patients. Each study includes one or more views (typically frontal and/or lateral) and a free-text report that often follows Findings and Impression sections. For data-efficient fine-tuning, we sample 1,000 studies per dataset (i.e. 1,000 from CheXpert-Plus and 1,000 from MIMIC-CXR) with a balanced ground-truth label set via CheXpert labeller [36]. We also exclude reports with missing sections. Following recent works [92], we report results on the test split and extract labels from both generated and GT reports by the CheXpert labeller. We further evaluate on the CheXpert Plus human gold set, whose expert-verified labels provide a more reliable basis for evaluation and stronger conclusions.
Baselines. Following the recent benchmarking study [92], we compare our proposed approach against a broad suite of methods: 28 task-specific RRG models spanning diverse architectures and 16 open-source LLMs/VLMs, as listed in Tables 1, 2 and 3. All task-specific algorithms are fine-tuned on at least the full training set of the dataset used in each experiment (CheXpert Plus or MIMIC-CXR), with some methods additionally incorporating multi-source corpora. Thus, our 1K training set sampled from CheXpert Plus, represents $\sim$0.45% of the standard 223K split; and only $\sim$0.079% of the 1.27M used by MambaXray [12], which is the current SOTA model. A similar ratio applies to MIMIC-CXR.
Evaluation Metric. We follow [92] to report clinical effectiveness as macro-averaged over labels: for each of the 14 conditions we compute Precision, Recall, and F1 on the test set, then take the unweighted mean across labels so frequent and rare conditions contribute equally.
Implementation Details.
We fine-tune a 3B Qwen-2.5-VL policy [4] on the 1,000 CheXpert Plus studies in an RL-only setting (no SFT/alignment), using Hugging Face TRL’s GRPOTrainer with transformers and accelerate. Training runs on four NVIDIA A10 GPUs (24 GB) with PyTorch 2.7.1/CUDA 12.1. We empirically set batch size $B=16$. For each prompt, we sample $K=8$ completions, compute the FactS reward per completion, and optimize the objective in Eq. (11). More detailed description of the hyper-parameters for OraPO can be found in the Supplementary.
4.2 Main Results and Analysis on CheXpert Plus
Against RRG task-specific algorithms.
As shown in Table 1, OraPO+FactS sets the new SOTA F1 = 0.341 (best) and Recall = 0.832 (best) on CheXpert Plus, significantly surpassing best-performing baselines by $\mathbf{+160.8\%}$ in recall. When compared with the most recent baseline, VLCI, our proposed method outperforms VLCI by $\mathbf{+109.2\%}$ in F1 and $\mathbf{+375.4\%}$ in recall. While our precision trails MambaXray-L, this is a deliberate trade-off between emphasizing per-finding coverage and accurate statements of detected findings. In practice, recall-oriented drafts are preferred. Since radiologists perform the final review, they tend to favour AI assistance that has higher sensitivity (even at the cost of precision) [31]. Overall, our method sets the new SOTA performance in F1 and recall, with three orders of magnitude less training data.
Against Alignment-based SFT for LLM/VLM. In this experiment, we compare our framework against R2GenGPT [94], a popular alignment-based SFT approach that is widely used for training large LLM/VLM. We report the performance of R2GenGPT [94] with different LLM/VLM backbones in Table 2. Our approach is trained with a compact 3B policy on 1K CheXpert Plus samples. Under this small-data setup, OraPO attains $\mathbf{F1=0.341}$ and $\mathbf{Recall=0.832}$, significantly outperforming larger aligned backbones trained on 223K. These results validate that our proposed OraPO with fact-level supervision is a superior learning paradigm to standard vision–language alignment, achieving higher performance with less than half backbone parameters and markedly less training data.
Evaluation on Human Gold Label Set. To obtain a more faithful evaluation, we assess models on the CheXpert validation set with human gold labels: 200 studies from 200 patients independently annotated by three board-certified radiologists, where a majority vote defines the ground truth [36]. Table 5 reports macro Precision/Recall/F1 across the 14 findings. With only 1K training cases, our method reaches 0.234/0.641/0.288, surpassing MambaXray-B, MambaXray-L and CheXagent in both Recall and F1. These gains arrive while using 0.1% of MambaXray’s data and around 0.01% of CheXagent’s.
We also evaluate two commercial APIs (GPT-4.1 [70] and GPT-5 Thinking [71]) on the CheXpert validation set with human gold labels, using the identical image input and report-generation prompt. GPT-5 is configured with a medium reasoning level. As reported in Table 5, our model outperforms GPT-4.1 by +110.2% in recall and +13.8% in F1. Relative to GPT-5 Thinking, our model achieves +52.3% higher recall, though F1 is -4.3% lower. Critically, compared to GPT-5 Thinking, our compact 3B model delivers real-time inference (3.3 s/image vs. 25.2 s/image), a lightweight and transparent footprint (3B vs. trillions of parameters), and cheaper deployment that eliminates per-token API fees. These advantages make our method substantially more cost-efficient and suitable for time-critical clinical use than calling commercial APIs. Overall, the results confirm that our improvements persist when measured against expert gold labels, while being dramatically more data-efficient.
4.3 Main Results and Analysis on MIMIC-CXRs
On MIMIC-CXR, our approach attains F1 = 0.357 and Recall = 0.891 using just 1K training samples, as shown in Table 3. This yields a +5.0% F1 gain over the best baseline (0.340, MambaXray-L) and a +153.8% recall gain over the strongest baseline recall (0.351, CXRMate). The markedly higher recall under macro averaging indicates substantially fewer omissions across both common and rare pathologies, aligning with clinical preference for high sensitivity and omission avoidance [62, 78, 76, 116, 69]. At the same time, the overall F1 is the best among all baselines, despite a precision of 0.242. Together with the CheXpert-Plus results, these findings show that OraPO generalises across datasets and achieves state-of-the-art clinical effectiveness with 2–3 orders-of-magnitude less data than prior approaches.
4.4 Ablation Study
To study the effectiveness and efficiency of our proposed method, we evaluate four model variants on CheXpert Plus using macro Precision, Recall, and F1. As reported in Table 4, the second row is a naïve GRPO-based RLHF baseline that uses accuracy as reward (as in DeepSeek-R1 [29]), which yields only marginal gains over the base model without any domain-specific training (row 1). Adding FactS to GRPO (row 3) lifts F1 from 0.089 to 0.291 $(+227\%)$, Recall from 0.162 to 0.605 (+274%), and Precision from 0.026 to 0.204 (+685%), showing that fact-level entailment provides dense, clinically aligned supervision. Layering OraPO on top (row 6) further improves F1 to 0.341 (+17.2% vs. FactS-only), with Recall improved by (+37.5%) and Precision improved by (+16.2%). Notably, OraPO+FactS with only 400 studies (row 4) attains F1 == 0.296, Precision == 0.217 and Recall == 0.732—–surpassing the FactS-only 1K setting (F1 == 0.291, Precision == 0.204, and Recall == 0.605)–—providing evidence of OraPO’s superb data efficiency. We further observe that the GRPO+SFT variant (with FactS) collapses to very low recall (0.176) and F1 (0.106), a -70.9% and -63.6% drop vs. FactS+GRPO, respectively. The core issue here is that plain SFT only imitates ground-truth prose: the model learns “what to say when it’s right,” but not “what not to say,” so reports that omit all true positives aren’t explicitly discouraged, leading to recall and F1 collapses. Beyond that, SFT on small datasets is known to be unstable and sample-inefficient: fine-tuning with limited data often over-fits, degrading generalisation [73, 22]. OraPO fixes this by explicitly pushing probability mass away from “no-TP” outputs such that the recall is effectively boosted when trained with limited set. Overall, FactS supplies faithful, sentence-level rewards, while OraPO converts zero-reward groups into preference updates, yielding higher precision and recall and faster convergence on scarce yet clinically important signals.
5 Conclusion
In this work, we present OraPO, an oracle-educated RL training algorithm for high-quality radiology report generation under tight data and compute budgets. To our knowledge, we are the first to couple DPO with GRPO: when GRPO exploration fails, we recycle those rollouts as negatives and apply a lightweight DPO step to recover useful gradients. In parallel, we design the FactS reward to provide dense, fact-level, clinically faithful supervision tailored to RRG. Comprehensive experiments show that OraPO sets the new SOTA performance while using 2–3 orders of magnitude fewer training samples. In future work, we target four directions: (1) Extend OraPO beyond RRG to sparse-signal tasks (e.g. multi-step math, code) where vanilla GRPO struggles. (2) Study scaling across model size and data/training budgets beyond our current 3B/1K setup. (3) Tackle non-verifiable rewards via rubric-driven signals (e.g. Rubrics as Rewards [28]). (4) Enrich FactS with fine-grained attribute entailment (location, laterality, size/severity, temporal change) to reduce diagnostic attribute misalignment.
Supplementary Material of OraPO: Oracle-educated Reinforcement Learning
for Data-efficient and Factual Radiology Report Generation
This supplementary material includes:
  • Algorithm 1: A detailed, step-by-step description of the proposed method.
  • Section A: Comprehensive implementation details.
  • Section B: Additional experimental results and analysis on the MIMIC-CXR dataset.
  • Section C: Qualitative studies, including examples of generated reports and assessments on extracted facts.
Appendix A More Implementation Details
[…]
Table 6: Hyperparameter settings of the proposed method (selected in underline).
Parameter Description Search Range / Setting
$B$ Effective Batch Size $\{\underline{16}\}$
LR Learning Rate $\{1e-6,\underline{2.5e-7}\}$
$K$ GRPO sampled group size $\{4,\underline{8},16\}$
$\alpha$ EMA momentum for zero-reward rate (ZRR) $\{0.4,\underline{0.5},0.6\}$
$w_{\min}$ Minimum DPO mixing weight $\{0.02,\underline{0.05},0.1\}$
$w_{\max}$ Maximum DPO mixing weight $\{\underline{0.15},0.3\}$
$\gamma$ Sharpening exponent for mapping ZRR to $w_{i}^{(t)}$ $\{1.0,\underline{2.0}\}$
Table 7: Experimental results (micro averaging) on the MIMIC-CXR dataset [8].
Algorithm Venue Precision, Recall, F1 Train Size
MedRAT [32] ECCV24 0.285, 0.265, 0.227 223K
MET [93] CVPR23 0.364, 0.309, 0.311 223K
KiUT [35] CVPR23 0.371, 0.318, 0.321 369K
DCL [46] CVPR23 0.471, 0.352, 0.373 223K
RGRG [83] CVPR23 0.524, 0.474, 0.498 223K
CoFE [47] ECCV24 0.489, 0.370, 0.405 223K
COMG [27] WACV24 0.424, 0.291, 0.345 223K
MPO [99] AAAI25 0.436, 0.376, 0.353 223K
MAN [80] AAAI24 0.411, 0.398, 0.389 223K
Med-LLM [55] MM24 0.412, 0.373, 0.395 223K
B-LLM [52] AAAI24 0.465, 0.482, 0.473 223K
EKAGen [6] CVPR24 0.517, 0.483, 0.499 223K
MambaXray-L [92] CVPR25 0.561, 0.460, 0.505 1.27M
MLRG [53] CVPR25 0.549, 0.468, 0.505 240K
Ours 0.342, 0.811, 0.481 1K
Algorithm 1 The proposed OraPO with FastS Reward for Radiology Report Generation
Input initial policy model $\pi_{\theta}^{\text{init}}$; dataset $\mathcal{D}$ with studies $(x,y^{\star})$ and prompts $p$; label set $L$; group size $K$; hyperparameters $\varepsilon,\lambda_{\mathrm{KL}},\tau,\alpha,\gamma,w_{\min},w_{\max},\beta,\xi$; outer loops $I$, inner steps $M$.
Output $\pi_{\theta}$
1:policy model $\pi_{\theta}\leftarrow\pi_{\theta}^{\text{init}}$
2:for $\text{iteration}=1,\ldots,I$ do
3:  reference model $\pi_{\mathrm{ref}}\leftarrow\pi_{\theta}$
4:  for $\text{step}=1,\ldots,M$ do
5:   Sample a minibatch $\{(x_{i},p_{i},y_{i}^{\star},z_{i}^{\star})\}^{B}\subset\mathcal{D}$
6:   Update the behaviour policy $\pi_{\theta_{\text{old}}}\leftarrow\pi_{\theta}$
7:   Sample $K$ reports $\{\hat{y}_{i,j}\}_{j=1}^{K}\sim\pi_{\theta_{\text{old}}}(\cdot\mid x_{i},p_{i})$ for each input radiology image $x_{i}$ in the batch
8:   Compute FactS reward: $\{r(x_{i},\hat{y}_{i,j})\}_{j=1}^{K}$ for each sampled report $\hat{y}_{i,j}$;
9:   for $\text{OraPO iterations}=1,\ldots,\mu$ do
10:     Compute GRPO objective: $L_{\mathrm{GRPO}}(x_{i},p_{i})$;
11:     Compute DPO objective: $L_{\mathrm{DPO}}(x_{i},p_{i})$;
12:     Compute OraPO objective: $\mathcal{L}_{\mathrm{OraPO}}$ by mixing $L_{\mathrm{DPO}}(x_{i},p_{i})$ and $L_{\mathrm{GRPO}}(x_{i},p_{i})$;
13:     Update the policy model $\pi_{\theta}$ by minimizing $L_{\mathrm{OraPO}}(x_{i},p_{i})$.
14:   end for
15:  end for
16:end for
Appendix B More experimental results and analysis on MIMIC-CXR
Table 7 complements our main macro-averaged results by reporting micro-averaged Precision, Recall, and F1 on MIMIC-CXR. Macro averaging, used in the main paper, treats each disease label equally and highlights performance on rare findings. Micro averaging instead weights labels by their frequency, answering a different question: across all positive label instances in the dataset, how many do we miss. This view is especially focusing on common conditions which dominate the label distribution.
Under this micro-averaging metric, OraPO remains highly competitive while using 2–3 orders of magnitude less data. With only 1 K training reports, OraPO attains a micro recall of 0.811 and micro F1 of 0.481. The best baseline recall is EKAGen at 0.483, so OraPO improves micro recall by 67.9% relative to this strong model. Clinically, such high recall means that, across all truly abnormal cases, our method misses far fewer findings, which is critical because false negatives on common diseases are usually more harmful than extra false positives.
Although our method achieves the highest recall, its micro F1 remains close to that of the strongest fully supervised approaches. The best baseline F1 is 0.505 (MambaXray-L and MLRG), whereas OraPO reaches 0.481, only a gap of 0.024. This trade-off is achieved with extreme data efficiency: compared with methods trained on the full 223K MIMIC-CXR split, OraPO uses 99.6% fewer labeled reports, and compared with MambaXray-L’s 1.27M training pairs, about 99.9% fewer. Overall, the micro results show that OraPO delivers clinically preferred high recall with competitive F1, while training with only 1K samples.
Appendix C Qualitative Study
In Figure 3, we present three examples that demonstrate how OraPO with FactS reward generates clinically accurate reports while maintaining factual grounding through atomic fact extraction and entailment checking.
Example 1 showcases ideal performance on a challenging case with cardiomegaly and edema. The model produces a clinically coherent narrative that captures the typical imaging pattern of cardiac-related pulmonary edema. Generated facts such as “cardiac silhouette is enlarged” and “diffuse interstitial opacities seen centrally” directly map to the ground-truth labels, with no false positives.
Example 2 highlights OraPO’s handling of complex cases with five ground-truth labels. The generated report correctly identifies all target pathologies, including the challenging bilateral nodular pulmonary lesions. While the model predicts pleural effusion absent from the ground-truth labels, the generated facts (“pleural effusions present” with “decreased interval changes”) suggest a false positive predicted by the model.
Example 3 illustrates robust multi-pathology detection spanning four conditions: lung opacity, consolidation, pneumonia, and pleural effusion. The model generates anatomically precise descriptions (“left perihilar opacity with patchy distribution”) that provide localization detail often absent in baseline systems. All ground-truth labels receive explicit supporting facts, demonstrating the FactS reward’s effectiveness in maintaining high recall across co-occurring pathologies. The model correctly attributes findings to specific lung regions, which aids radiologist review.
Across these examples, OraPO consistently produces factually grounded reports where each diagnostic statement traces back to verifiable atomic facts that entail specific labels. This stands in contrast to fluency-optimized baselines that generate plausible prose without guaranteed factual alignment. The FactS reward’s sentence-level supervision enables the model to maintain clinical accuracy even when handling rare pathologies or multi-label complexity, directly addressing the recall-precision trade-off critical in clinical deployment.
Here’s a concise qualitative analysis paragraph you can place right below Fig. 3:
C.1 Qualitative NLG Metric Analysis
Although the ROUGE-L F1 values in Figure 3 are modest, they largely reflect surface mismatch rather than clinical errors: OraPO is not trained to imitate the phrasing or template of GT reports. Instead of imitating the GT writing style, our method directly optimizes a fact-level reward. The model first extracts atomic clinical statements, and then checks whether these statements are entailed by the ground-truth labels. This training objective naturally prioritizes clinical correctness and verifiability, not surface-level similarity. In contrast, reference-overlap metrics such as ROUGE mainly measure wording overlap. They reward matched phrasing but do not penalize factual mistakes, missing findings, or even cross-sentence contradictions in long reports. Concretely, our FactScore-based reward treats the report as its own rationale, maps sentences to label-level predictions, and yields dense, interpretable feedback to the policy. In the three qualitative cases, the generated statements are clinically grounded and aligned with the GT labels: Example 1 captures cardiomegaly and edema with supporting facts; Example 2 correctly identifies all target pathologies (with one flagged false-positive effusion); and Example 3 localizes a left perihilar opacity and supports consolidation/pneumonia/effusion—all consistent with the label set despite low lexical overlap with GT wording. This behavior is consistent with our quantitative results. OraPO is deliberately recall-oriented, which is clinically the safer operating point. On CheXpert Plus, it achieves state-of-the-art label-level recall and F1 while using only about 0.1% of the usual training data, and it also attains very high micro-recall. Therefore, low n-gram overlap with the reference should not be interpreted as low clinical faithfulness.
Image GT Reports GT Labels Our Generated Reports Predicted Labels and Supporting Facts ROUGE-L F1
Refer to caption Single frontal view of the chest demonstrates indistinct pulmonary vasculature bilaterally, likely representing mild pulmonary edema. Mild cardiomegaly is seen, unchanged from prior exam. No soft tissue or bony abnormalities are seen. Cardiomegaly, Edema Ribs: The thoracic rib contour is smooth and complete. There is mild opacity seen in the mid to left hemithorax. Pectoral Girdle: Both clavicles are seen in their normal positions. Diaphragm: The left diaphragm appears increased in curvature without significant deviation. Aortic Arch/Trunk: The aortic arch appears well defined. Cardiomediastinum: There are diffuse interstitial opacities seen centrally on the image, consistent with pulmonary edema. The right border of the heart is poorly defined and the cardiac silhouette is enlarged. The pulmonary vascularity is increased centrally but diffuse through the lung fields and the cardiac silhouette appears to be expanding to the right of the sternum. Pulmonary Vasculature: The pulmonary vasculature appears increased centrally. Lung Opacity: There is diffuse sub-pleural opacity centrally, with extension towards the left hemithorax. Increased pulmonary vasculature seen centrally, consistent with pulmonary edema. No other areas of consolidation seen. This appearance is consistent with pulmonary edema. There are small sub-pleural opacities seen centrally. - Cardiomegaly: ’Cardiac silhouette is enlarged’ and ’Cardiac silhouette appears to be expanding to the right of the sternum’ both indicate cardiomegaly. - Edema: ’Diffuse interstitial opacities seen centrally’, ’Pulmonary vascularity is increased centrally’, ’Pulmonary vascularity is diffuse through the lung fields’, and ’Appearance is consistent with pulmonary edema’ all indicate edema.” 0.102
Refer to caption Within the right middle lobe there is a large consolidation, consistent with pneumonia, which obscures the right heart border. A small left upper lobe calcified granuloma is present. There is mild calcification of the aortic arch. No pleural effusions. Mild cardiomegaly. No bony abnormalities are appreciated. Cardiomegaly, Lung Opacity, Lung Lesion, Consolidation, Pneumonia Cardiomediastinal silhouette: mildly increased in size; cardiac contour is flattened Nodular pulmonary lesions present bilaterally - no growth or interval increase Pleural effusions present - decreased interval changes Pneumonic opacities present, right mid and left lobe, bilateral, but limited in extent Right diaphragmatic elevation - Cardiomegaly: ’Cardiomediastinal silhouette is mildly increased in size’ directly supports presence of cardiomegaly. - Lung Opacity: Facts ’Pneumonic opacities present in right mid lobe’, ’Pneumonic opacities present in left lobe’, and ’Nodular pulmonary lesions present bilaterally’ all describe opacities. - Lung Lesion: ’Nodular pulmonary lesions present bilaterally’ directly supports presence of lung lesions. - Consolidation: Facts ’Pneumonic opacities present in right mid lobe’, ’Pneumonic opacities present in left lobe’, and ’Pneumonic opacities are bilateral’ support presence of consolidation. - Pneumonia: Facts ’Pneumonic opacities present in right mid lobe’, ’Pneumonic opacities present in left lobe’, and ’Pneumonic opacities are bilateral’ support presence of pneumonia. - Pleural Effusion: ’Pleural effusions present’ directly supports presence of pleural effusion. 0.114
Refer to caption Initial chest radiograph at 1130 hours demonstrates left lower lobe focal air space opacity which could represent pneumonia. The remainder of the lungs are clear. The cardiomediastinal silhouette is normal. There is blunting of the left costophrenic angle which could represent a small pleural effusion. Follow-up radiograph 1530 hours demonstrates no interval change. Lung Opacity, Consolidation, Pneumonia, Pleural Effusion Single ventricular image. Cardiac silhouette is no grossly enlarged (LVED: 55-57 mm, LVED: 55-57 mm) and changes relative to previous films. No evidence of marked cardiomegaly in the presence of mild respiratory straining. Left perihilar opacity with patchy distribution. Small pleural effusion present, mild volume and interval increase noted.Cardiac silhouette interval increase noted. - Lung Opacity: Fact ’Left perihilar opacity with patchy distribution’ describes an opacity, confirming existence. - Consolidation: Fact ’Left perihilar opacity with patchy distribution’ describes an opacity consistent with consolidation. - Pneumonia: Fact ’Left perihilar opacity with patchy distribution’ describes an opacity that could represent pneumonia. - Pleural Effusion: Facts ’Small pleural effusion present’, ’Mild volume increase in pleural effusion’, and ’Interval increase in pleural effusion’ all confirm pleural effusion. 0.164
Figure 3: X-ray image and its corresponding ground-truth, along with the output of our model generation report on the ChexPert Plus dataset. The mismatch sentence in the reports are highlighted using different colors.